Jane has $12:00 in quarters and dimes. She has five times as many dimes,d, as quarters.q. Determine the number of quarters she has.
Answer by stanbon(75887)

Jane has $12:00 in quarters and dimes. She has five times as many dimes,d, as quarters.q. Determine the number of quarters she has.
====
Quantity: d = 5q
Value:: 10d + 25q = 1200 cents
----
Substitute for "d" and solve for "q":
10(5q) + 25q = 1200
75q = 1200
q = 16 (# of quarters)
----
Solve for "d":
d = 5q
d = 5*16 = 80 (# of dimes)
==============================
Cheers,
Stan H.
=============